CR =6%, i = 5.4%, 4-year bond, bond price?
Answer #1

Information provided:

Par value= future value=$1,000

Time= 4 years

Coupon rate= 6%

Coupon payment= 0.06*1,000= $60

Interest rate= 5.4%

The price of the bond is computed by calculating the present value.

Enter the below in a financial calculator to compute the present value:

FV= 1,000

N= 4

PMT= 60

I/Y= 5.4

Press the CPT key and PV to compute the present value.

The value obtained is 1,021.0795

Therefore, the price of the bond is $1,021.08.

In case of any query, kindly comment on the solution.